### v1

# Efficient Parallel Implementation of the Multiplicative Weight Update Method for Graph-based Linear Programs

Abstract.
[…]
In this work, we implement and test an efficient parallel algorithm for solving positive LP relaxations, and apply it to graph problems such as densest subgraph, bipartite matching, vertex cover and dominating set. We accelerate the algorithm via a new step size search heuristic. Our implementation uses sparse linear algebra optimization techniques such as fusion of vector operations and use of sparse format. Furthermore, we devise an implicit representation for graph incidence constraints. We demonstrate the parallel scalability with the use of threading OpenMP and MPI on the Stampede2 supercomputer. We compare this implementation with exact libraries and specialized libraries for the above problems in order to evaluate MWU’s practical standing for both accuracy and performance among other methods. Our results show this implementation is faster than general purpose LP solvers (IBM CPLEX, Gurobi) in all of our experiments, and in some instances, outperforms state-of-the-art specialized parallel graph algorithms.
copyright: none
1. Introduction
Designing general graph libraries for massively-parallel machines is challenging as combinatorial graph algorithms often have limited concurrency and arithmetic intensity (Lumsdaine et al., 2007). Many have parallel depth at least proportional to graph diameter, though in practice these algorithms contain sufficient concurrency to achieve high-efficiency on shared-memory machines (Nguyen et al., 2013; Shun and Blelloch, 2013; Shun et al., 2015; Dhulipala et al., 2019). Designing efficient distributed-memory parallelizations is more challenging, though it has been achieved with use of graph partitioning (Sui et al., 2010), formulations via sparse matrix products (Buluç and Gilbert, 2011; Buluç et al., 2009; Solomonik et al., 2017).
Due to this nature, combinatorial algorithms may not scale well on large graphs or large diameter graphs. In addition, some parallelization techniques do not easily generalize between combinatorial algorithms so they don’t naturally allow code reuse. Finally, for certain graph problems, like densest subgraph  (Charikar, 2000), a combinatorial approach does not lend algorithms that are efficient, simple, and almost exact.
Another method for solving graph problems is to reformulate or relax (Williamson and Shmoys, 2011; Vazirani, 2013) the graph problem into a linear program (LP). Recent theoretical breakthroughs in both sequential and parallel algorithms for several graph problems (Andoni et al., 2020; Li, 2020; van den Brand et al., 2020a; van den Brand et al., 2020b; Sherman, 2013; Liu and Sidford, 2020) indicate that carefully designed LP solvers can enable algorithms with tunable accuracy that have lower cost and depth than their combinatorial counterparts. Given a matrix $\bm{A}\in\mathbf{R}^{m\times n}$, vector $\bm{c}\in\mathbf{R}^{n}$, and vector $\bm{b}\in\mathbf{R}^{m}$, an LP is the optimization problem,
(1) $\max_{\bm{x}\in\mathbf{R}^{n}}\langle\bm{c},\bm{x}\rangle\text{ s.t. }\bm{Ax}\leq\bm{b},$
where $\langle\bm{c},\bm{x}\rangle=c_{1}x_{1}+c_{2}x_{2}+\ldots+c_{n}x_{n}$, and $\bm{u}\leq\bm{v}$ means $u_{i}\leq v_{i}\ \forall i$. (Williamson and Shmoys, 2011; Vazirani, 2013).
Typically, to solve a general LP, one employs variants of the simplex or interior point methods. While these algorithms are efficient in the sequential setting, they often have limited parallelism. LP solving is P-Complete, which implies that a poly-logarithmic depth parallel algorithm is believed unlikely to exist (Karp and Ramachandran, 1989).
However, many graph problems can be solved via LPs that contain only positive entries, and these class of LPs admit efficient parallelizable solvers if some approximation is allowed. These approaches yield a parallel algorithm with depth that is a polynomial in $1/\epsilon$ and $\log n$ (the number of variables), and is independent of the structure of the constraint matrices $\bm{P}$ and $\bm{C}$. It outputs a result that is a $(1+\epsilon)$-approximation, which, for a maximization problem entails a relative error of $(1-\epsilon)$ in the objective value, and for a minimization problem, a relative error of $(1+\epsilon)$.
The polynomial scaling with $1/\epsilon$ has been improved in recent work. Our implementation is based on an accelerated version of a more recent algorithm with an improved depth of $\tilde{O}(\epsilon^{-3})$ for mixed problems and $\tilde{O}(\epsilon^{-2})$ for pure covering or pure packing problems. The only other previously published study  (Makari et al., 2013) of a distributed and parallel LP solver built using this approach implemented and compared runtimes of two earlier algorithms: MPCSolver and Young’s algorithm, with $\tilde{O}(\epsilon^{-5})$ and $\tilde{O}(\epsilon^{-4})$ depth, respectively. Further details on theoretical developments in approximately solving positive LPs in parallel are provided in Section 2.
The algorithm we focus on is from Mahoney et. al. (Mahoney et al., 2016), which offers the fastest theoretical performance for pure packing, pure covering, and mixed packing and covering LPs in a parallel setting. We describe these problems and their associated graph problems formally in Section 3. This algorithm we call MWU since it utilizes the weights vector from the more general MWU method (Arora et al., 2012). Despite the algorithm’s strong theoretical foundation, little is known about its empirical performance, especially for solving real-world graph problems over large datasets. Furthermore, understanding how this algorithm compares to other LP solvers and specialized graph libraries is an open question.
In this paper, we create a practical and scalable implementation of a parallel (1+$\epsilon$)-approximate positive LP solver by adapting the MWU algorithm with the current fastest theoretical performance, and we also provide a comprehensive empirical study comparing MWU against exact solvers, specialized libraries, and previous related work. We also test a line search method that finds the largest step-size permitted without violating theoretical guarantees. This step-size search reduces in the number of overall MWU iterations in practice by multiple orders of magnitude. Our results suggest that MWU can be fast in both theory and practice, with runtimes comparable to even specialized libraries but with the added benefit of, one, being generalizable and, two, more versatile, as it can benefit from high-performance sparse linear algebra libraries. We believe this marks MWU as a viable alternative for solving graph problems approximately on a large dataset. Overall, our paper makes the following contributions:
  • the first shared and distributed-memory implementation of a general-purpose approximate solver based on Mahoney et. al. (Mahoney et al., 2016) for positive LPs, called MWU
  • a step size search method for MWU that empirically reduces the iteration count by up to three orders of magnitude without adding significant overhead
  • an efficient implementation of MWU using an implicit representation of common constraint matrices observed in graph LPs as well as other standard techniques for SLA. We show these optimizations provide good scalability and lead up to 5.2x speedup relative to MWU implemented with PETSc (a library for parallel sparse linear algebra).
  • the first comparison of an approximate positive LP solver against general LP solvers as well as specialized parallel graph algorithms. In particular, MWU finds $(1+\epsilon)$-relative ($\epsilon=0.1$) solutions up to 3-2800x and 5-1180x faster than CPLEX and Gurobi (which find exact solutions for implicit ILPs and exact fractional solutions for relaxed LPs), respectively, for solving several graph problems on large real-world graphs on the Stampede2 supercomputer using KNL compute nodes.
2. Background on Linear Program Solvers
2.1. Positive LP Solvers
Fast approximate solvers for positive LPs in the sequential settings have been developed since the early 1990’s (Plotkin et al., 1995; Grigoriadis and Khachiyan, 1994), and there is extensive and continued attention to this line of work. We mostly focus on parallel algorithms and refer the reader to (Quanrud, 2019; Wang, 2017) for extensive pointers. Luby and Nisan provide the first parallel algorithm for explicit positive LPs which obtain a $(1+\epsilon)$-approximation in $\tilde{O}(\epsilon^{-4})$ iterations [^1] We write $\tilde{O}(f(n))$ to be proportion to $f(n)$ and a polylogarithmic of $f(n)$, i.e., $\tilde{O}(f(n))\propto O(f(n)\log^{O(1)}(f(n)))$ for pure packing and covering LPs (Luby and Nisan, 1993). Young clarified and extended this work to solve the more general mixed packing and covering LPs in parallel in $\tilde{O}(\epsilon^{-4})$ iterations (Young, 2001). The dependence on $\epsilon$ has remained unchanged for over 10 years until the work of Allen-Zhu and Orecchia, who solve pure packing and pure covering LPs in parallel with $\tilde{O}(\epsilon^{-3})$ iterations (Allen-Zhu and Orecchia, 2016). They also obtained better dependence on $\epsilon$ in the sequential setting (Allen-Zhu and Orecchia, 2019) for pure packing and covering LPs (see also (Mahoney et al., 2016)).
Recently, Mahoney et. al. (Mahoney et al., 2016) utilized ideas of Young (Young, 2014) on faster near-linear time sequential solver to develop new parallel algorithms for positive LPs. For mixed packing and covering LPs, their algorithm converges in $O\big{(}\log(m_{p}+m_{c})\log(n/\epsilon)/\epsilon^{3}\big{)}$ iterations and is also work-efficient, i.e., work is near linear to the number of nonzeros in the LP. For pure packing and pure covering LPs, the dependence on $\epsilon$ is $\epsilon^{-2}$ instead. When $\epsilon$ is not too small, these algorithms have low depth in the PRAM model.
Empirical studies of these fast approximate algorithms have been mainly limited to relatively small problems. Koufogiannakis and Young adapt a sequential mixed packing and covering LP and show it outperforms simplex on randomly generated binary matrices with dimension up to $2$ (Koufogiannakis and Young, 2014). Allen-Zhu and Orecchia compare their $\tilde{O}(\epsilon^{-1})$-dependent sequential algorithm (Allen-Zhu and Orecchia, 2014) to the algorithms of Luby and Nisan (Luby and Nisan, 1993) and Awerbuch and Khandekar (Awerbuch and Khandekar, 2009) for solving pure packing LPs with a randomly generated matrix of size $60\times 40$. Jelic et. al. implement a parallel primal-dual method on GPUs to solve positive LPs, although their constraint matrices are randomly generated binary matrices with dimensions up to 25000 (Jelić et al., 2015). The most closely related work to ours is that of Makari et. al. (Makari et al., 2013), who implement a gradient descent algorithm to solve generalized matching on large real-world and synthetic graphs. Their implementation, based on the algorithm from (Awerbuch and Khandekar, 2009), has a $\tilde{O}(1/\epsilon^{5})$ number of iterations, and they confine their attention to a single graph problem. In this paper we focus on only obtaining fractional solutions to the LP problems we consider. The work in  (Makari et al., 2013) also has a rounding step to convert the fractional solution to an integral matching.
2.2. The MWU Algorithm
We now introduce MWU, the algorithm of Mahoney et. al. (Mahoney et al., 2016) for approximately solving the standard mixed packing and covering LP in parallel. This section does not contain our modifications for improving the empirical performance, which are found in Section  4.
First, we start describing how MWU solves the mixed packing and covering feasibility LP  (Young, 2001; Mahoney et al., 2016). We will discuss how to modify this into an optimization problem later in the section. The feasibility LP is:
(2) $\exists\ {\bm{x}\in\mathbf{R}^{n}}\text{ s.t. }\bm{Px}\leq\mathbb{1},\bm{Cx}\geq\mathbb{1},\bm{x}\geq\mathbb{0},$
where $\bm{P}$ and $\bm{C}$ are nonnegative. Vectors $\mathbb{1}$ and $\mathbb{0}$ are the all ones and zeros vector, respectively. The algorithms we consider seek a $(1+\epsilon)$-relative approximation: that is, a solution $\bm{x}$ such that $\bm{Px}\leq(1+\epsilon)\mathbb{1}$ and $\bm{Cx}\geq\mathbb{1}$.
MWU (Algorithm 1) ensures that both packing and convering constraints, $\bm{Px}\leq\mathbb{1}$ and $\bm{Cx}\geq\mathbb{1}$, are approximately satisfied by approximating $\max(\bm{Px})$ and $\min(\bm{Cx})$ with smoothed maximum and minimum functions,
$\mathrm{smax}_{\eta}(\bm{x})$ $=\frac{1}{\eta}\log\Big{(}\sum\limits_{i=1}^{n}\mathrm{exp}(\eta\cdot x_{i})\Big{)},$
$\mathrm{smin}_{\eta}(\bm{x})$ $=-\frac{1}{\eta}\log\Big{(}\sum\limits_{i=1}^{n}\mathrm{exp}(-\eta\cdot x_{i})\Big{)},$
where $\eta>2$ is a smoothing parameter. The MWU algorithm and step size search make use of their gradients,
$\nabla\mathrm{smax}_{\eta}(\bm{x})$ $=\frac{\mathrm{exp}(\eta\cdot\bm{x})}{\langle\mathbb{1},\mathrm{exp}(\eta\cdot\bm{x})\rangle},\nabla\mathrm{smin}_{\eta}(\bm{x})=\frac{\mathrm{exp}(-\eta\cdot\bm{x})}{\langle\mathbb{1},\mathrm{exp}(-\eta\cdot\bm{x})\rangle}.$
For more details on these functions, see Chapter 2 of (Quanrud, 2019).
Algorithm 1 Multi-Update MWU Method for Mixed Packing and Covering LPs
1:procedure MWU($\bm{P}\in\mathbf{R}_{+}^{m_{P}\times n}$, $\bm{C}\in\mathbf{R}_{+}^{m_{C}\times n}$, $\epsilon$ )
2:    $\eta\leftarrow 10\log(m)/\epsilon$ where $m:=m_{P}+m_{C}$
3:    $x_{i}\leftarrow\frac{\epsilon}{n\|\bm{P}_{:,i}\|_{\infty}}\ \forall i\in[n]$triangleright $\|\bm{x}\|_{\infty}=\max_{i}|x_{i}|$
4:    while constraints not approximately satisfied and $\bm{C}\neq\varnothing$ do
5:        $\bm{g}\leftarrow\bm{P}^{\mathsf{T}}\nabla\mathrm{smax}_{\eta}(\bm{Px})$
6:        $\bm{h}\leftarrow\bm{C}^{\mathsf{T}}\nabla\mathrm{smin}_{\eta}(\bm{Cx})$
7:        $d_{i}\leftarrow\frac{1}{2\eta}\mathrm{max}\{0,1-\frac{g_{i}}{h_{i}}\}\cdot x_{i}\ \forall i$
8:        if $\max(\bm{d})=0$ then
9:           Return “INFEASIBLE”
10:        $\bm{x}\leftarrow\bm{x}+\bm{d}$
11:        $\bm{C}\leftarrow\{c_{i}\ :\ c_{i}^{T}x<1\}$triangleright Keep unsatisfied constraints
12:    return $x$
The algorithm initializes the vector $\bm{x}$ with small values so that each starting packing constraint is at most $\epsilon$ (Line 3). The smoothing parameter $\eta$ is set so that both $\mathrm{smax}_{\eta}$ and $\mathrm{smin}_{\eta}$ are within an $\epsilon$ additive error of max and min, respectively. In each MWU iteration, the algorithm multiplicatively updates $\bm{x}$. This is done by defining a step or update vector, $\bm{d}$, where $d_{i}$ is a multiple of $x_{i}$ (Line 7), and adding $\bm{d}$ to $\bm{x}$ (Line 10). Vectors $\bm{g}$ and $\bm{h}$, which are gradients of the smoothed max packing and min covering constraints, respectively, are also utilized to define $\bm{d}$ (Lines 56). In particular, $\bm{d}$ is an approximate solution to the Lagrangian relaxation,
(3) $\begin{split}\exists\bm{d}\in\mathbf{R}^{n}_{\geq 0}\text{ s.t. }&\langle\bm{w}_{p},\bm{Pd}\rangle=\langle\bm{P}^{T}\bm{w}_{p},\bm{d}\rangle\leq 1,\\ &\langle\bm{w}_{c},\bm{Cd}\rangle=\langle\bm{C}^{T}\bm{w}_{c},\bm{d}\rangle\geq 1,\end{split}$
where $\bm{w}_{p}=\nabla\mathrm{smax}_{\eta}(\bm{P}x)$ and $\bm{w}_{c}=\nabla\mathrm{smin}_{\eta}(\bm{C}x)$.
If the positive LP is infeasible, then there exists some MWU iterations where $\bm{d}=\mathbb{0}$, in which case the algorithm reports the LP is infeasible (Line 8(Mahoney et al., 2016). Assuming otherwise, the theoretical analysis guarantees MWU will return an $(1+\epsilon)$-relative solution. Throughout the algorithm, we drop satisfied covering constraints, as these can unnecessarily slow down progress (Line 11). Finally, the algorithm returns $\bm{x}$ when all the covering constraints are satisfied or when there exists no covering constraints.
We note that Algorithm 1 can also solve pure packing or pure covering LPs, which are, respectively,
$\max\ \langle\mathbb{1},\bm{x}\rangle\text{ s.t. }\bm{Px}\leq\mathbb{1},\bm{x}\geq\mathbb{0},\bm{x}\in\mathbf{R}^{n}$
$\min\ \langle\mathbb{1},\bm{x}\rangle\text{ s.t. }\bm{Cx}\geq\mathbb{1},\bm{x}\geq\mathbb{0},\bm{x}\in\mathbf{R}^{n}.$
For example, to solve a pure packing LP, we embed the objective function as the added constraint, $\frac{1}{M}\mathbb{1}^{T}\bm{x}\geq 1$, where $M$ is the estimate of the maximum value. Then we do binary search over $M$, using Algorithm 1 to determine if the resulting mixed packing and covering LP is a feasible. Since there is one covering constraint, then $\mathrm{smin}_{\eta}(\bm{Cx})=\min(\bm{Cx})$. This exact approximation permits one to scale the step direction (Line 7) by a factor of $2$ in the theoretical analysis, which improves the number of iterations by a factor of $\epsilon$ (Mahoney et al., 2016). Also, noting $\bm{C}=\frac{1}{M}\mathbb{1}^{T}$ and $\nabla\mathrm{smin}_{\eta}(\bm{Cx})=1$, we have $\bm{h}=\frac{1}{M}\mathbb{1}$ (Line 6), so we do not need to explicitly compute $\bm{h}$. Solving a pure covering LP is done via a similar transformation. Solving a mixed covering and packing optimization problem, likewise, involves embedding the constraint that corresponds to the direction of optimization.
3. Graph Problems as Positive LPs
We now consider several graph problems and LP formulations for them. We first define integer programming (IP) formulations which exactly model the underlying graph problem. We then obtain an LP by relaxing the integrality constraints. Note that for some problems, such as bipartite matching and densest subgraph, the solution to the LP relaxation matches the IP’s solution (i.e., the solution is integral), whereas for NP-hard problems dominating set and vertex cover, there is an integrality gap between the solution to the LP relaxation and IP. We refer the reader (Vazirani, 2013; Williamson and Shmoys, 2011) for the role of LP relaxations in the development of approximation algorithms, and also  (Li et al., 2020) for a performance study on recovering an integral output from a relaxed solution for dominating set as a case study. In this paper, we do not consider rounding for LPs with integrality gaps in both our implementation and in performance comparisons.
Let $G=(V,E)$ be an unweighted, undirected graph where $V$ is the set of vertices and $E$ is the set of edges, with $n=|V|$ and $m=|E|$. For simplicity, we assume $G$ has no self-loops. Note that our formulations can be extended towards weighted graphs as well. For a vertex $v\in V$, let $N(v)$ be the neighbor vertices of $v$ ($v$ is not included in $N(v)$), and $\mathrm{inc}(v)$ be the set of edges incident to $v$.
The neighbor relations between the vertices of $G$ can be represented as an adjacency matrix, $\bm{A}\in\{0,1\}^{n\times n}$, which is symmetric and has a nonzero for each edge $e\in E$. The incidence relation between the vertices and the edges can be represented as a vertex-edge incidence matrix, $\bm{M}$, where
(4) $\bm{M}_{u,e}=\begin{cases}1\ :\ u\in e,e\in E\\ 0\ :\ \text{otherwise}\end{cases},\ \bm{M}\in\{0,1\}^{n\times m}.$
In this work, we consider four graph applications: maximum matching, dominating set, vertex cover, and densest subgraph.
Maximum Matching (match). A matching is a subset $F\subset E$ of edges such that no vertex is incident to more than one edge in $F$. We can write this optimization problem as the IP,
(5) $\begin{split}\max\sum\limits_{e\in E}x_{e}\ \text{s.t.}&\ \sum\limits_{e\in\mathrm{inc}(v)}x_{e}\leq 1\ \forall v\in V\\ &\ x_{e}\in\{0,1\}\ \forall e\in E.\end{split}$
The $x_{e}$ variables indicate whether edge $e$ is in set $F$. The constraints are defined over the vertices of the graph such that at most one edge ($x_{e}$) in the incident edges of a vertex $v$ ($\mathrm{inc}(v)$) can be selected for matching set $F$. When the input graph is a bipartite graph, we call this problem maximum bipartite matching (bmatch).
(a) match
(b) dom-set
(c) vcover
(d) dense-sub
Figure 1. Four graph problems run on the same graph. Variables of LP with a nonzero value are highlighted in red. An example of matching is given in Figure 1a. The set of edges in the matching are marked with thick red lines.
Given the vertex-edge incidence matrix $\bm{M}$ of the graph, we can write the LP relaxation for maximum matching as the pure packing LP,
(6) $\max\ \langle\mathbb{1},\bm{x}\rangle\text{ s.t. }\bm{Mx}\leq\mathbb{1},\bm{x}\geq\mathbb{0},\ \bm{x}\in\mathbf{R}^{m}.$
It is well-known that this LP relaxation has no integrality gap for bmatch while for general graphs there is an integrality gap of $2/3$; there is an exponential sized exact LP relaxation for general graph matching but we do not consider it here.
Dominating Set (dom-set). Dominating set is the problem of finding the smallest subset of vertices $S\subseteq V$ such that every vertex in the graph is either in $S$ or is a neighbor of a vertex in $S$. We can formulate the dom-set problem as the IP,
(7) $\begin{split}\min\ \sum\limits_{v\in V}x_{v}\ \text{ s.t. }&\ x_{v}+\sum\limits_{u\in N(v)}x_{u}\geq 1,\ \forall v\in V\\ &\ x_{v}\in\{0,1\}\ \forall v\in V.\end{split}$
The variable $x_{v}$ indicates whether vertex $v$ is in set $S$ or not. The constraints are defined over the vertices such that either vertex $v$ itself or one of its neighbors is in the set $S$. The LP relaxation is a pure covering LP,
(8) $\begin{split}\min\ \langle\mathbb{1},\bm{x}\rangle\text{ s.t. }(\bm{I}+\bm{A})\bm{x}\geq\mathbb{1},\bm{x}\geq\mathbb{0},\ \bm{x}\in\mathbf{R}^{n},\end{split}$
where $\bm{I}$ is the identity matrix.
Vertex Cover (vcover). In the vertex cover problem the goal is to find the smallest subset of vertices $S\subseteq V$ such that every edge has one of its endpoints in $S$ (hence $S$ covers all the edges). A simple IP formulation is
(9) $\begin{split}\min\ \sum\limits_{v\in V}x_{v}\ \text{s.t.}&\ x_{u}+x_{v}\geq 1\ \forall(u,v)\in E\\ &\ x_{v}\in\{0,1\}\ \forall v\in V.\end{split}$
Here, the $x_{v}$ variables determines whether a vertex $v$ is in set $S$. These variables are defined over the vertices. There is one constraint per edge. The LP relaxation is a pure covering LP,
(10) $\begin{split}\min\ \langle\mathbb{1},\bm{x}\rangle\text{ s.t. }\bm{M}^{\mathsf{T}}x\bm{\geq}\mathbb{1},\bm{x}\geq\mathbb{0},\bm{x}\in\mathbf{R}^{n},\end{split}$
where $\bm{M}^{T}$ is the transpose of the vertex-edge incidence matrix of the graph.
Densest Subgraph (dense-sub). Densest subgraph finds a subgraph $S\subset G$ that maximizes the edge to vertex count ratio, i.e., $|E(S)|/|S|$. The LP, as formulated in (Charikar, 2000), is
(11) $\begin{split}\max\ \sum\limits_{e\in E}x_{e}\ \text{s.t.}&\ x_{e}\leq y_{u},x_{e}\leq y_{v}\ \ \forall e=(u,v)\in E\\ &\ \sum\limits_{v\in V}y_{v}\leq 1\\ &\ x_{e},y_{v}\geq 0\ \forall v\in V,\forall e\in E.\end{split}$
The variables $x_{e}$ represent the edges and the variables $y_{v}$ represent the vertices, which are no longer binary. Since this problem is not a positive LP, we consider its dual (Boob et al., 2020),
(12) $\begin{split}\min D\text{ s.t.}&\ z_{u,e}+z_{v,e}\geq 1\ \forall e=(u,v)\in E\\ &\ \sum\limits_{e\in\text{inc}(v)}z_{v,e}\leq D\ \forall v\in V\\ &\ z_{v,e}\geq 0\ \forall v\in V,e\in\mathrm{inc}(v).\end{split}$
While (12) is still not a positive LP, we can convert it to a mixed packing and covering LP by fixing $D$ to be a constant and treating (12) as a feasibility problem instead. We find an approximate minimum value to (12) via binary search for $D$ and solving the feasibility LP for each choice.
Unlike previous LPs, the variables $z_{v,e}$ represent a vertex-edge pair instead of vertices or edges, so we require new constraint matrices. Let $\bm{I}$ be an identity matrix of size $m$, and let the function $\mathrm{interweave}$ take two equally-sized matrices and put the first and second matrices’ first columns as the first two columns of the combined matrix, then their second columns as the next pair of columns, and so on. We call the resulting matrix $\bm{W}$ the interweaved identity matrix, where
(13) $\bm{W}_{e,2e},\bm{W}_{e,2e+1}=1,\ \forall e\in E,\bm{W}\in\{0,1\}^{n\times 2m}$
In order to model the vertex-edge pair variables, we can form a matrix called vertex-edge pairs matrix. Vertex-edge pairs matrix will have a column for each vertex $v$ and edge ($u,v$). Specifically, we can form the vertex-edge pair matrix, $\bm{O}\in\{0,1\}^{n\times 2m}$, from a graph $G$ as follows:
(14) $\begin{split}\bm{O}_{u,2e+b}=\begin{cases}1&:b=0,e=(u,v)\in E\\ 1&:b=1,e=(v,u)\in E\\ 0&:\text{otherwise}\end{cases}\end{split}$
Then the feasibility variant of the dual to densest subgraph is
(15) $\exists\bm{z}\in\mathbb{R}^{2m}\text{ s.t.}\ \bm{W}z\geq\mathbb{1},\ \bm{O}\bm{z}\leq D\cdot\mathbb{1},\ \bm{z}\geq\mathbb{0}.$
4. MWU with Line Search
We now introduce algorithmic modifications to Algorithm 1 that improve its performance in practice, which we review in 7.3. We write it in Algorithm 2 and mark the modifications in red. Similar to Algorithm 1, we can adapt this algorithm to solve pure packing (e.g., maximum matching) and pure covering (e.g., dominating set, vertex cover) LPs more efficiently, as described in Section 2.2.
Algorithm 2 Multi-Update MWU Method with step size search for Mixed Packing and Covering LPs
1:procedure MWU_STEPSIZE_SEARCH($\bm{P}\in\mathbf{R}_{+}^{m_{P}\times n}$, $\bm{C}\in\mathbf{R}_{+}^{m_{C}\times n}$, $\epsilon$ )
2:    Initialize $\eta$, and $x_{i}$ as in Algorithm 1
3:    $\bm{y}\leftarrow\bm{Px}$, $\bm{z}\leftarrow\bm{Cx}$
4:    while constraints not approximately satisfied and $\bm{C}\neq\varnothing$ do
5:        $\bm{g}\leftarrow\bm{P}^{\mathsf{T}}\nabla\mathrm{smax}_{\eta}(\bm{y})$triangleright Packing gradient
6:        $\bm{h}\leftarrow\bm{C}^{\mathsf{T}}\nabla\mathrm{smin}_{\eta}(\bm{z})$triangleright Covering gradient
7:        $d_{i}\leftarrow\frac{1}{2\eta}\mathrm{max}\{0,1-\frac{g_{i}}{h_{i}}\}\cdot x_{i}\ \forall i$triangleright New step direction
8:        if $\max(\bm{d})=0$ then
9:           Return “INFEASIBLE”
10:        $\bm{d}^{(y)}\leftarrow\bm{Pd}$, $\bm{d}^{(z)}\leftarrow\bm{Cd}$
11:         $\alpha\leftarrow\operatorname{StepSize}(\bm{d},\bm{y},\bm{z},\bm{d}^{(y)},\bm{d}^{(z)},\eta)$triangleright Step size search
12:        if $\alpha<1$ then
13:           Return “INFEASIBLE”
14:        $\bm{x}\leftarrow\bm{x}+{\color[rgb]{1,0,0}\alpha}\cdot\bm{d}$
15:        $\bm{y}\leftarrow\bm{y}+{\color[rgb]{1,0,0}\alpha}\cdot\bm{d}^{(y)}$, $\bm{z}\leftarrow\bm{z}+{\color[rgb]{1,0,0}\alpha}\cdot\bm{d}^{(z)}$
16:        $\bm{C}\leftarrow\{c_{i}\ :\ c_{i}^{T}x<1\}$
17:    return $x$
We store the packing and covering constraints $\bm{y}=\bm{Px}$ and $\bm{z}=\bm{Cx}$ as well as $\bm{d}^{(y)}=\bm{Pd}$ and $\bm{d}^{(z)}=\bm{Cd}$ (Line 3 and 10) to minimize the number of sparse matrix-vector products, or SpMVs. The step direction $\bm{d}$ is unchanged (Line 7). While the algorithm drops satisfied constraints (Line 16), in practice we keep satisfied constraints since this simplifies the implementation and we did not find it impacts the convergence on the problems we tested.
The sub-routine StepSize (Line 11) takes the step vector $\bm{d}$ and constraint vectors, and returns a step size $\alpha>0$. We call this modification step size search and design algorithms for it in the next subsection. When $\alpha<1$, we report that finding a solution is infeasible, because otherwise a step size of $\alpha=1$ is always possible due to the theoretical analysis of Mahoney et. al. (Mahoney et al., 2016). Therefore, we call a step size $\alpha=1$ found without step size search the standard step size. Assuming $\alpha\geq 1$, we scale the step direction $\bm{d}$ by $\alpha$ and add it to $\bm{x}$ (Line 14). Afterwards, we update $\bm{y}=\bm{Px}$ and $\bm{z}=\bm{Cx}$ without SpMVs (Line 15). Note that $\alpha$ may be large enough so that $\bm{Px}=\bm{y}+\alpha\cdot\bm{d}^{(z)}\geq 1$, in which case we terminate MWU.
4.1. Line Search as a Constrained Optimization Problem
In this section, we consider algorithms for finding a step size for StepSize. When selecting $\alpha$, we want it to be sufficiently large to accelerate MWU convergence while ensuring that we recover a feasible solution to (2).
First, we consider the case of the mixed packing covering LP. One of the qualities of Mahoney et. al’s algorithm is that the algorithm reaches a $(1+\epsilon)$-approximation when the difference in a potential function $f(x)=\frac{1}{\eta}(\mathrm{smax}_{\eta}(\bm{Px})-\mathrm{smin}_{\eta}(\bm{Cx}))$ becomes sufficiently small (Mahoney et al., 2016). Each step that their algorithm takes is non-increasing on $f(x)$.
We can show that in order for the potential function to be non-increasing, $f(x^{(t+1)})-f(x^{(t)})=\Psi(\alpha)-\Phi(\alpha)\leq 0$ where,
$\Phi(\alpha)$ $=\mathrm{smin}_{\eta}(\bm{C}(\bm{x}+\alpha\cdot\bm{d}))-\mathrm{smin}_{\eta}(\bm{Cx})$
$\Psi(\alpha)$ $=\mathrm{smax}_{\eta}(\bm{P}(\bm{x}+\alpha\cdot\bm{d}))-\mathrm{smax}_{\eta}(\bm{Px}).$
This gives us an equivalent invariant $f(\alpha)=\Phi(\alpha)/\Psi(\alpha)\geq 1$. Therefore, if we find a step size $\alpha$ for which this invariant holds, then for this $\alpha$ we can still say $f(x)$ is non-increasing, and furthermore, if we can reach a point where $\bm{Px}\leq(1+\epsilon)\mathbb{1}$ and $\bm{Cx}\geq\mathbb{1}$ then we have converged to a feasible solution. Hence, our method is to find the largest step size $\alpha>0$ such that the “bang-for-buck” value is at least one, or
(16) $f(\alpha)=\Phi(\alpha)/\Psi(\alpha)\geq 1,$
For pure packing and pure covering problems we instead have these invariants, respectively:
(17) $\begin{split}\langle\mathbb{1},\alpha\bm{d}\rangle/\Psi(\alpha)\geq 1\\ \langle\mathbb{1},\alpha\bm{d}\rangle/\Phi(\alpha)\leq 1\\ \end{split}$
We now show MWU with line search maintains the same theoretical properties as MWU with the standard step size (Mahoney et al., 2016). Recall $\bm{x}\in\mathbf{R}^{n}$ and $m$ is the number of rows in the matrices $\bm{P}$ and $\bm{C}$.
Theorem 4.1 ().
MWU with line search (Algorithm 2) either returns an $(1+\epsilon)$-relative approximate solution, i.e., an $x\geq\mathbb{0}$ such that $\bm{Px}\leq(1+\epsilon)\mathbb{1}$ and $\bm{Cx}\geq\mathbb{1}$, or correctly reports the LP is infeasible. The number of iterations is at most $\tilde{O}(\epsilon^{-3})$, where $\tilde{O}$ hides polylogarthmic dependence on $n$, $m$, and $\epsilon$.
The proof is similar to the one shown in (Mahoney et al., 2016), which implicitly sets the step size to $\alpha=1$. There will be two main differences in the convergence proof, which we highlight here. First, the proof of correctness in  (Mahoney et al., 2016) shows the potential function, defined as $\mathrm{smax}_{\eta}(\bm{Px})-\mathrm{smin}_{\eta}(\bm{Cx})$, is monotonically decreasing by taking a first-order approximation of smooth max and min. On the other hand, our bang-for-buck invariance (16) explicitly ensures this monotonicity property. Second, the argument in (Mahoney et al., 2016) upper bounds the number of MWU iterations by lower bounding the values in step direction vector $\bm{d}$. Since line search only increases $\bm{d}$ because $\alpha\cdot\bm{d}\geq\bm{d}$, then line search can only decrease the number of MWU iterations.
While line search can decrease the number of iterations (in fact, quite significantly in our experiments), finding a step size increases the work per iteration. In the following lemma, we leverage the monotonicitiy of $f(\alpha)$ to design efficient line search algorithms.
Proposition 4.2 ().
$f$ is monotonically decreasing for $\alpha\in\mathbf{R}_{+}$.
Proof.
We show that as $\alpha$ increases $\Psi(\alpha)/\alpha$ is increasing while $\Phi(\alpha)/\alpha$ is decreasing, hence
$f(\alpha)=(\Phi(\alpha)/\alpha)/(\Psi(\alpha)/\alpha)$
is decreasing. Note that $\Psi$ is convex since $\mathrm{smax}_{\eta}$ is convex, and $\Phi$ is concave since $\mathrm{smin}_{\eta}$ is concave. Since $\Psi$ is convex, $\Psi(\alpha)\leq\Psi(0)+\alpha\Psi^{\prime}(\alpha)=\alpha\Psi^{\prime}(\alpha)$. Hence, we can show that $\Psi(\alpha)/\alpha$ is increasing, since
$(\Psi(\alpha)/\alpha)^{\prime}=\frac{1}{\alpha}(\Psi^{\prime}(\alpha)-\Psi(\alpha)/\alpha)\geq 0.$
Analogously, since $\Phi$ is concave, the inequalities above are reversed, and so $\Phi(\alpha)/\alpha$ must be strictly decreasing in $\alpha$. ∎
4.2. Implementing Line Search
To approximate the maximum step size $\alpha^{*}$ satisfying (16), we first perform exponential search to find an integer $p$ where $f(2^{p})\geq 1$ and $f(2^{p+1})<1$. By Proposition 4.2, $\alpha^{*}\in[2^{p},2^{p+1})$. Next, we run binary search starting with a lower and upper bound of $l=2^{p}$ and $u=2^{p+1}$, and update the lower and upper bounds so that $l$ (resp. $u$) is the largest (smallest) value such that $f(l)\geq 1$ ($f(u)<1$). Once we find an $\epsilon$-relative step size, or when $\frac{u-l}{l}\leq\epsilon$, we return $l$. We formalize the aforementioned procedure in Algorithm 3.
When $f(\alpha)\geq 1$ and $\max\big{(}\bm{C}(\bm{x}+\alpha\cdot\bm{d})\big{)}\geq 1$ (Line 4), this means the step size can lead MWU to completion, so we immediately return $\alpha$. Moreover, binary search makes use of $\bm{y}=\bm{Px}$, $\bm{z}=\bm{Cx}$, $\bm{d}^{(y)}=\bm{Pd}$, and $\bm{d}^{(z)}=\bm{Cd}$, where $\bm{x}$ is our current solution and $\bm{d}$ is the computed MWU update direction, to avoid additional SpMVs.
Algorithm 3 Finding a step size via binary search
1:procedure BinSearch($\{\bm{y},\bm{d}^{(y)}\}\in\mathbf{R}^{m_{p}}_{\geq 0}$, $\{\bm{z},\bm{d}^{(z)}\}\in\mathbf{R}^{m_{c}}_{\geq 0}$, $\epsilon$)
2:    $\alpha\leftarrow 1$
3:    while $f(\alpha)\geq 1$ do $\triangleright$ Exponential search. See (16)
4:        if  $\mathrm{min}(\bm{z}+\alpha\cdot\bm{d}^{(z)})\geq 1$  then
5:           return $\alpha$triangleright Return early if constraints are satisfied
6:        $\alpha\leftarrow 2\cdot\alpha$
7:    $lb,ub\leftarrow\alpha/2,\alpha$
8:    while $ub-lb>(1-\epsilon)lb$ do $\triangleright$ Binary search
9:        $\beta\leftarrow\mathrm{avg}(lb,ub)$
10:        if $f(\beta)\geq 1$ then $\triangleright$ See (16)
11:           $lb\leftarrow\beta$
12:        else
13:           $ub\leftarrow\beta$
14:    $\alpha\leftarrow lb\text{ and }\textbf{return}{}\ \alpha$
We can derive another line search method by using Newton’s method, which has the update,
$\alpha_{k+1}=\alpha_{k}-g(\alpha_{k})/g^{\prime}(\alpha_{k}),\text{ where }g(\alpha)=f(\alpha)-1.$
Because Newton’s method converges when its solution is in the neighborhood of the optimal solution, we require estimates of $\alpha^{*}$ to ensure convergence. We do so via a warm start for Newton’s search, where we set our initial $\alpha_{0}$ to the previous optimal step size, if available, or use exponential search. The reason for the former strategy is we observed in our tests that the optimal step size between two MWU iteration are relatively close. Finally, we note that once Newton’s method converges to some solution, it may not strictly satisfy (16). Thus, we multiplicatively decrease the solution by a factor of $(1-\epsilon)^{p}$ for some integer $p$ ($p$ is typically small) until (16) is satisfied.
5. Software Optimizations and Parallelization
We now describe the details of our implementation and parallelization of linear algebra operations within MWU. To efficiently perform sparse matrix-vector products with matrices introduced in Section 3, such as the vertex-edge adjacency matrix, we leverage implicit representations derived from a standard sparse vertex-vertex adjacency matrix data structure. We accelerate vector operations with loop fusion and vectorization.
5.1. Shared-Memory Optimizations
We design implicit SpMVs for matrices that arise in graph-based positive LPs such as vertex-edge incidence matrices. We adapt previous fusion techniques for the needs of MWU framework (Roth and Kennedy, 1998).
5.1.1. Choice of Matrix Format
To efficiently traverse the non-zeros in the adjacency matrix during SpMVs, we use the Compressed Sparse Blocks (CSB) format (Buluç et al., 2009), which can achieve good cache locality for both SpMVs of the matrix and its transpose.
CSB divides the matrix into two-dimensional $r\times k$ tiles. Each tile is represented as a list of tuples, where each tuple stores the non-zeros in column major order in coordinate (COO) format. The group of tiles that belong to $r$ consecutive rows is called a row-block while the group of tiles that belong to $c$ consecutive columns is called a column-block. In our implementation, we store the tiles in row-major order.
Figure 2. CSB representation and its SpMV operation.
Similar to (Aktulga et al., 2014), we parallelize $\bm{y}=\bm{A}\bm{x}$ and $\bm{y}=\bm{A}^{T}\bm{x}$ over the row-blocks and column-blocks, respectively. Provided the tile size ($r\times k$) is selected carefully, the block of the input vector ($\bm{x}$) processed by a tile and output vector $\bm{y}$ corresponding to a row-block (updated by a single thread) is contained in private L1 or L2 caches. Figure 2 illustrates the parallelization of SpMV using CSB format.
5.1.2. Implicit Representations
In each iteration of MWU, we perform one SpMV with the constraint matrix and with its transpose. As seen in Section 3, the constraint matrices of many graph-based LPs are the vertex-edge incidence matrix of the graph. We notice that the edge $(u,e)$ encoded in an incidence matrix can be described over over a vertex and a vertex pair $(u,(v,w))$ where the pair $(v,w)$ represents an edge in the adjacency matrix. In addition, in the incidence matrix, the value is $1$ if $u=v$ or $u=w$ and 00 otherwise. Therefore, we can store the nonzero data in an incidence matrix implicitly as an adjacency matrix in memory. This reduces the memory cost of the constraint matrix by about half and therefore also reduce the number of accesses to the memory subsystem, particularly cache.
Having an implicit representation means that linear algebra operations on these matrices can also be expressed implicitly by formulating the computations on the adjacency matrix, $\bm{A}$.
We emulate SpMVs for $\bm{M}$ and $\bm{O}$ by storing the edge, row, and column index for each non-zero in $\bm{A}$, which we denote by $e$, $r$, and $c$, respectively. For example, we can compute $\bm{y}=\bm{M}\bm{x}$ by accumulating element $x_{e}$ to $y_{r}$ and $y_{c}$ for each $(r,c,e)$ in $\bm{A}$, and likewise for $\bm{y}=\bm{O}\bm{x}$, then evaluating $y=y_{r}+y_{c}$. To parallelize this SpMV while avoiding race conditions, we first traverse in row major order while reading the row indices of $\bm{A}$, then in column major order while reading the column indices.
We compute $\bm{y}=\bm{M}^{T}\bm{x}$ by accumulating elements $x_{r}$ and $x_{c}$ to $y_{e}$ for all $(r,c,e)$ in $\bm{A}$, and likewise for $\bm{y}=\bm{O}^{T}\bm{x}$. Parallelizing these SpMVs are straightforward, since we accumulate to any given element of $\bm{y}$ two and one times, respectively.
5.1.3. Loop Fusion and Vectorization Opportunities
In each iteration of the MWU algorithm, we do several vector operations and our augmentation to step size search adds many vector operations, too. For some problems, such as vertex-cover or densest subgraph, the vectors in these operations have size $|E|$, which means they can be as costly as a single SpMV.
Since these vector operations loop to apply simple arithmetic to each element in the vector, combining multiple vector operations in one pass via loop fusions can accelerate these methods. We identify two operations for fusion: (1) the gradient calculations using smax and smin (Lines 5 and 6) and (2) the calculation of the new step direction  (16). In both cases, loop fusion can reduce memory accesses and facilitate automatic vectorization.
5.2. Distributed Parallelization
Distributed-memory parallelization of vector operations and explicit sparse matrix vector products in MWU can be done with standard techniques. Therefore, in this section, we will focus on describing and analyzing the benefits of the implicit representation for distributed-memory communication.
We use the same implicit representation described in Section 5.1.2. We leverage a 2D matrix distribution of the adjacency matrix to perform implicit SpMVs with the incidence matrix. A 2D data layout is communication-efficient for matrix vector products since each processor computes on only $n/\sqrt{p}$ entries and contributes to $n/\sqrt{p}$ outputs (for an $n\times n$ matrix on a $\sqrt{p}\times\sqrt{p}$ processor grid). They are commonly employed for parallel processing of adjacency matrices (Buluç and Gilbert, 2011).
With a 2D layout of the adjacency matrix, we perform vertex-edge incidence products with twice the communication cost of an adjacency matrix product. To do so, we store vector information corresponding to edges in the same processor layout as the adjacency matrix $A$. This means that for an edge $(u,v)$ in $A$, the machine owning the edge would store vector information for indices corresponding to $u$ and to $v$.
For simplicity, we assume a square processor grid. With this approach, the product with the vertex-edge incidence matrix, $\bm{y}=\bm{M}\bm{x}$, requires only a reduction of contributions to $\bm{y}$ along rows and columns of the processor grid. While for the product $\bm{y}=\bm{M}^{T}\bm{x}$, only a broadcast of entries of $\bm{x}$ along rows and columns of the processor grid is needed. In both cases, each processor sends or receives a subvector of size $O(\sqrt{n}/{p})$.
6. Experimental Setup
6.1. System Setup
We use Intel Knights Landing (KNL) nodes on the Stampede2 supercomputer as our testbed. Each KNL node has 68 1.4 GHz cores. Each core has a 32 KB private L1 cache, and 2 neighboring cores share a 1 MB L2 cache. KNL processors also support AVX2 and AVX-512 vector instructions. Each Stampede2 node has 112 GB of memory capacity with 96 GB DRAM and 16 GB MCDRAM used in cache mode.
6.2. Implementations
MWU Implementations. We implement two different versions of the MWU framework presented in Algorithm 2: (1) MWU-PETSc, and (2) MWU-opt. MWU-PETSc relies on an efficient parallel BLAS library PETSc (Balay et al., 2019) while MWU-opt is our hand-optimized implementation using optimizations discussed in Section 5. One modification in our implementation of Algorithm 2 is we do not drop satisfied constraints (i.e., we skip Line 16). This simplifies the implementation, and we did not find this affects convergence to the optimal solution.
PETSc (Balay et al., 2019) is an efficient distributed sparse BLAS library with a Python interface (petsc4py). On Stampede2, we use PETSc with MKL version 19.1.1. We use C++ for our MWU-opt optimized implementation and OpenMP for parallelism. We compile our code with Intel compiler version 19.1.1 and enable -O3, and -mAVX512 compiler flags.
We run the MWU-opt implementation by binding threads to physical cores using numactl –physcpubind. For MWU-PETSc, we find that creating $N$ processes each with 1 thread gives the best performance.
MWU has two parameters: $\epsilon$ and max_iter. $\epsilon$ controls the accuracy of the solution, while max_iter controls the maximum number of MWU iterations, where an LP requiring more than max_iters MWU iterations is deemed infeasible. We set $\epsilon=0.1$ and max_iter=5000. The choice of max_iter is ad-hoc and based on observations that MWU finds $(1+\epsilon)$-relative solutions in at most 1000 iterations across most LP problems and inputs. To verify correctness, we compare the solution from MWU with an exact solution, if available, and report instances where the relative error is greater than $\epsilon$. We implement all applications discussed in Section 3 with MWU.
General LP Solvers. We compare our optimized MWU implementation to general LP solvers. We use IBM CPLEX (Cplex, 2009) and Gurobi (Gurobi Optimization, 2021), both in multi-threaded settings. If the problem is an ILP, we do not round the fractional solution.
For CPLEX, we set the run mode to opportunistic to achieve the fastest (but non-deterministic) runtime. For Gurobi, we use the concurrent optimization setting, which concurrently runs primal simplex, dual simplex, and the barrier method. We report the fastest runtime out of these three methods. When the barrier method finishes first, we report its runtime before crossover (unless otherwise noted) for a more fair comparison to MWU, which outputs fractional solutions. We implement all applications discussed in Section 3 with both CPLEX and Gurobi. Finally, we limit the solve time for both CPLEX and Gurobi to 4 hours. All other parameters were set to defaults.
Specialized Algorithms. We consider optimized custom implementations as baselines for the two implicitly integral LP problems: bipartite matching and densest subgraph. For the former, we use ms-bfs-graft (Azad et al., 2016), which employs the serial Karp-Singer greedy initialization step followed by a specialized breadth-first searches to find augmenting paths. For the latter problem, we used the Graph Based Benchmark Suite’s (Dhulipala et al., 2020) (GBBS) approximate densest subgraph algorithm, which implements Charikar’s greedy 2-approximation algorithm (Charikar, 2000). Both ms-bfs-graft and GBBS are implemented in C++ with OpenMP. We compile ms-bfs-graft using OpenMP and the Intel compiler (version 19.1.1) with the -O2 flag. We compiled GBBS with the g++ compiler version 9.1.0.
6.3. Input Graphs
We select a variety of real-world and synthetic undirected graphs from the SuiteSparse Matrix Collection (Davis and Hu, 2011) and list them in Table 1. Our real-world graphs come from diverse domains, such as a road, social, and user-product network. We also use two sets of synthetic graphs, the first set being random geometric graphs (rgg) which have a planar-like structure, and the second set being Kronecker graphs (kron) from Graph500 which show a strong community structure.
Note that none of the graphs we selected are bipartite, which is required in bmatch. To obtain bipartite graphs, we read the input adjacency matrix as a biadjacency matrix, meaning that the rows and columns of the matrix correspond to the left and right sets of vertices, respectively, where edges can only go between between vertices in different sets.
Table 1. List of real-world and synthetic graphs
Graphs (Abv.) $|V|$ $|E|$
usroads (usroads) 129,164 330,870
com-Amazon (amazon) 334,863 1,851,744
coPapersCiteseer (papers) 434,102 32,073,440
hollywood-2009 (hollyw) 1,139,905 113,891,327
com-Orkut (orkut) 3,072,441 234,370,166
kron-X X=$2^{17}$-$2^{21}$approxX×$\times 80$
rgg-Y Y=$2^{17}$-$2^{24}$approxY×$\times 15$
7. Experimental Results
In this section, first, we compare our implementation to state-of-the-art software including general LP solvers, CPLEX and Gurobi, to specialized parallel implementations for particular graph problems, and to the parallel implementation of another multiplicative weights update algorithm from Makari et. al  (Makari et al., 2013).
Then, we evaluate the effectiveness of our algorithmic improvements and software optimizations. We start by finding how the incorporation of a step size search reduces the number of MWU iterations. We then test the performance improvements from our software optimizations and its scalability by comparing performance between MWU-PETSc and MWU-opt.
7.1. Comparison of MWU to Other Algorithms
Table 2. Runtime (in seconds) of MWU-opt compared to other optimization libraries and custom applications. Cells with a dash indicate the algorithm took 4+ hours to run or had a memory error, with the exception of vcover with Gurobi on all the kron graphs and hollyw., which had a ConstraintError.
bmatch dom-set vcover dense-sub
MWU CPLEX Gurobi graft MWU CPLEX Gurobi MWU CPLEX Gurobi MWU CPLEX Gurobi gbbs
rgg-15 0.08 9.15 6.93 0.07 0.38 5.56 3.60 2.38 3.99 7.88 0.16 24.01 12.90 0.04
rgg-16 0.09 21.23 14.38 0.14 0.34 15.62 10.64 3.37 10.72 6.86 0.22 54.57 25.61 0.05
rgg-17 0.24 45.35 31.68 0.32 0.68 43.11 22.91 5.65 28.98 15.90 0.54 141.95 119.43 0.09
rgg-18 0.15 114.31 76.76 0.71 3.54 561.23 49.01 9.85 82.43 38.42 0.66 349.34 344.72 0.13
rgg-19 0.54 283.79 170.81 1.86 2.08 3,045.28 111.90 23.48 114.63 86.94 1.41 1,202.33 877.80 0.21
rgg-20 0.43 716.66 406.77 4.44 7.79 - 255.15 44.84 292.40 226.01 2.34 4,081.44 2,017.68 0.32
rgg-21 0.85 2,186.03 917.63 11.12 17.25 - 555.60 87.95 659.18 504.65 5.15 - - 0.56
rgg-22 2.81 - - 30.67 46.37 - 1,313.81 183.67 - - 13.21 - - 0.90
rgg-23 3.92 - - 80.99 247.68 - - 367.86 - - 22.40 - - 1.69
rgg-24 21.00 - - 226.12 115.02 - - 856.06 - - 74.87 - - 3.18
kron-16 3.80 95.61 136.6 0.00 1.38 19.22 23.22 4.53 81.83 - 1.81 3,169.42 - 0.11
kron-17 1.87 200.79 335.18 0.01 10.97 63.81 59.44 55.16 194.90 - 2.85 8,053.81 - 0.17
kron-18 2.60 462.66 642.27 0.01 3.07 214.97 155.06 33.63 414.28 - 10.53 - - 0.26
kron-19 4.38 - - 0.01 10.24 657.11 379.04 53.06 2,354.63 - 12.88 - - 0.43
kron-20 10.36 - - 0.02 32.51 2,292.96 1,048.8 82.93 3,091.02 - 24.30 - - 0.67
kron-21 32.33 - - 0.04 584.85 7,072.73 2,342.73 210.92 - - 56.28 - - 1.13
usroads 0.07 20.24 12.30 0.04 1.49 16.30 8.85 1.07 12.31 5.22 0.27 35.25 16.80 0.03
amazon 0.71 93.93 126.34 0.05 22.14 115.10 98.99 2.13 72.38 - 1.52 750.42 1,040.63 0.09
papers 6.44 3,549.09 542.19 0.33 10.14 35.14 49.15 149.95 767.30 443.05 3.78 - 3,945.37 0.39
hollyw. 39.13 - - 0.56 43.86 - - 130.50 - - 24.79 - - 1.89
orkut 29.99 - - 29.12 162.24 - - 334.70 - - 201.09 - - 3.18
We now compare the MWU-opt implementation of MWU with Newton’s method and all the software implementation optimizations to other state-of-the-art optimization libraries and custom implementations (ms-bfs-graft for bmatch and GBBS for dense-sub). All experiments are run with 64 threads on a single KNL node. Table 2 shows the execution times to find $(1+\epsilon)$-relative solutions for four positive LPs on various graphs where $\epsilon=0.1$. A “-” in a cell means that the input graph was either too large to be processed by the library, or the runtime exceeded 4 hours.
Comparison with Exact LP Solvers.
For all LP solvers, we do not do rounding as post-processing. Therefore, for the exact solvers, we have integral solutions to graphs problems that have implicitly integral LPs, and exact fractional solutions for the relaxed LPs with integrality gaps. For approximate solvers, we are not guaranteed an integral solution on integral LPs. Note that CPLEX and Gurobi return exact solutions for the target LPs, while MWU finds an $(1+\epsilon)$-relative solution with a target value of $\epsilon=0.1$. We find that our MWU-opt implementation is able to find an $\epsilon=0.1$ solution in all cases except bmatch problem with rgg-20. However, even for this case, our error rate is 0.104.
Our results show MWU-opt consistently outperforms CPLEX and Gurobi libraries. For bmatch, dom-set, vcover, and dense-sub graph LPs, MWU-opt is up to 2548x (rgg-21), 1482x (rgg-19), 43x (kron-19), and 2860x (kron-17) faster than CPLEX, respectively. Although Gurobi is generally faster than CPLEX, MWU-opt outperforms Gurobi by up to 1070x (rgg-21), 55x (rgg-19), 5x (rgg-21), and 816x (rgg-20) for bmatch, dom-set, vcover, and dense-sub graph LPs, respectively, before crossover occurs. When comparing the time when after crossover or one of the simplex methods from Gurobi terminates (whichever comes first), the relative speedups are 1462x (rgg-21), 3510x (rgg-19), 10x (usroads), and 878x (rgg-20) for bmatch, dom-set, vcover, and dense-sub graph LPs, respectively. In addition to significant speedups, we also observe that MWU is capable of running much larger problems. For instance, both CPLEX and Gurobi can only solve kron-21 for dom-set but not for dense-sub, the latter which contains three times more nonzeros than the former. Moreover, both LP solvers fail to solve any of the problems with the largest graphs, hollywood and orkut.
Comparison with Custom Implementations. We compare MWU-opt performance to ms-bfs-graft for bmatch and to GBBS for dense-sub. The MS-BFS algorithm returns an exact solution. MWU-opt returns an approximate solution with $\epsilon=0.1$. The MS-BFS algorithm  (Azad et al., 2016) initializes with a serial Karp-Singer greedy step and finds augmenting paths in parallel using specialized BFS. The performance of ms-bfs-graft heavily depends on the graph structure. In general, we observed the MWU-opt can outperform ms-bfs-graft for graphs with planar structures by 1.8-22.4x for the rgg graphs and usroads. On the other hand, for graphs which contain a strong community-structure or vertices with high degrees, ms-bfs-graft outperforms MWU. For example, amongst the kron graphs, ms-bfs-graft can be up to 450x faster than MWU-opt. For these problems, bmatch general spends much less time on the grafting process than with planar-structured graphs, which we often find to be the dominating cost of bmatch.
For dense-sub, GBBS implements Charikar’s greedy 2-approximation algorithm for densest subgraph (Charikar, 2000), but the relative error is usually much better in practice (but worse than $\epsilon=0.1$). Again, we run MWU-opt with $\epsilon=0.1$. We observe that GBBS always outperforms MWU, achieving a maximum speedup of 63.2x and minimum speedup of 4x.
Comparison with Previous Work. We compare MWU (Algorithm 2) and an implementation of a gradient descent algorithm with adaptive error (Makari et al., 2013), which uses a less theoretically efficient multiplicative weights update algorithm but is the only other distributed study of multiplicative weights update methods on graph problems. Their paper compared their algorithm, called MPCSolver, against their implementation of Young’s parallel algorithm for feasibility generalized matching, which is a mixed packing and covering LP, for an $(1+\epsilon)$-relative solution ($\epsilon=0.05$) and found that the implementation of MPCSolver outperformed Young’s algorithm. We provide a detailed description of the problem, datasets, and gradient descent algorithm in Appendix A.1.
We run the same experiment as them with MWU-opt, using the same datasets as well: the Netflix and KDD datasets (Bennett et al., 2007; Dror et al., 2012). Because both algorithms solve the same LP with a multiplicative weight update approach, there are only minor differences in vector operations between the two algorithms. Consequently, for this section, we compare iteration counts rather than time between the two algorithms. The two proposed algorithms are compared in Figure 3. For MWU, we consider both the standard step size and the Newton’s method for step size search. The gradient descent data is manually extracted from (Makari et al., 2013) using WebPlotDigiter (Rohatgi, 2020). The plot shows both MWU with Newton’s method and gradient descent with adaptive error find a $(1+\epsilon)$-relative solution in less than 2000 iterations, whereas MWU with standard step size converges much more slowly. Moreover, MWU with Newton’s method incurs 10×10$\times$10 × and 41×41$\times$41 × fewer iterations than gradient descent for Netflix and KDD, respectively.
These results highlight the effectiveness of using a step size strategy, such as Newton’s method, over the standard step size. Furthermore, MWU with Newton’s method converges more rapidly than gradient descent with an adaptive error. However, since both methods use heuristics to accelerate the method, testing additional positive LPs and datasets would be needed for a comprehensive understanding of the trade-offs between MWU and gradient descent. The heuristic that MPCSolver uses prematurely stops the algorithm once it detects that the per-iteration decrease in constraint violation falls below a threshold.
Figure 3. Max violation, defined as $\max\{0,\max(\bm{Px})-1,1-\min(\bm{Cx})\}$ for MPCSolver, which is a gradient descent algorithm with adaptive error (Makari et al., 2013), and MWU (Algorithm 2) with standard step size and Newton’s method.
7.2. Parallel Scalability of MWU
We now analyze the strong scaling behavior of MWU-PETSc and MWU-opt. Figure 4 displays the speedup with respect to single-threaded execution of the MWU-opt implementation.
When executing on a single node, all LP problem types are run along a range of thread counts from single threaded to 68 threads, which is the maximum number of hardware threads on one machine. Here, MWU-opt is able to achieve speedup over 16x with 68 threads in 90% of experiments and over 32x in 50% of experiments. Overall, the MWU-opt implementation achieves speedups of 13-55x on 68 threads compared to its single threaded runtimes. The largest differences between MWU-opt and MWU-PETSc are observed on graph applications where we use specialized matrices such as vcover, bmatch, and dense-sub problems. High parallel efficiency in MWU-opt is achieved due to load balancing and high locality in matrix-vector multiplications of transposed specialized matrices and vector operations. On the other hand, we see that the MWU-opt can only achieve 2-3x speedup compared to MWU-PETSc for dominating set LP (dom-set) application. Note that, for dom-set, MWU-opt can only make use of format selection and memory access minimization optimizations for SpMV operations.
For multi-node results, we execute MWU-opt and MWU-PETSc with 64 MPI processes per node and 1 thread per process (this was the most performant configuration for MWU-PETSc on a single machine). We only run experiments where the total number of processes is square, as that is a requirement for our implicit representation. We run all algorithms with Newton’s optimization for step size search and limit Newton’s methods to 5000 iterations. On distributed memory, except for vertex cover on rgg-24, MWU-opt runs faster than MWU-PETSc at scale. For the distributed problem, we also observe almost linear scaling for all graphs except rgg-24 in MWU-opt. A matrix-vector product on the incidence matrix of banded matrices, like rgg-24, corresponds to a 1D problem, so a 1D parallelization (e.g., row-wise distribution of the matrix) is more efficient than a 2D distribution, which is the layout we use.
For MWU-PETSc, we observe good scaling on dom-set, which we expect as the LP for this problem does not use implicit representation and is a 1D problem. We observe poor scaling in all other problems on all graphs except rgg-24. For rgg-24, MWU-PETSc achieves good performance due to its internal representation (Balay et al., 2019), which, we believe communicates only the vector entries needed by each processor based on the sparsity pattern of rows assigned to it. However, MWU-PETSc performs extremely poorly on dense-sub when we use multiple processors and does not complete in under 2 hours. In conclusion, for general graphs, the implicit 2D representation scales well compared to a explicit 1D representation.
(a) dom-set orkut
(b) dom-set hollyw.
(c) match hollyw.
(d) match kron-21
(e) bmatch hollyw.
(f) bmatch kron-21
(g) vcover orkut
(h) vcover rgg-24
(i) dense-sub orkut
(j) dense-sub rgg-24
Figure 4. Scalability of MWU-opt and MWU-PETSc. Axes are in $log_{2}$ scale. All values are normalized to single-threaded execution of MWU-opt. We omit the results for MWU-PETSc, if the execution time is slower than single-threaded execution of MWU-opt.
7.3. Improvements from Step Size Strategy
We first evaluate the effectiveness of step size search (Section 4). We run MWU-PETSc using 64 MPI proccesses, each with 1 thread and list the results for rgg-18 in Table 3. We choose this graph since we have exact solutions for all five graph problems, and the runtime with standard step size is not too large. The speedups for other graphs are within an order of magnitude of the ones listed here.
Table 3. Convergence of MWU to find $(1+\epsilon)$-relative ($\epsilon=0.1$) solution on rgg-18 with standard step strategy (Std), binary search (Bin), and Newton’s method (Nwt). For the latter two step strategies, we use the previous step size as the initial step size.
# MWU iters
Avg # step
size iters
Time (sec)
Std Bin Nwt Bin Nwt Std Bin Nwt
match 25477 13 13 8.31 4.86 79.3 0.87 0.94
bmatch 28210 15 13 8.00 5.07 261 1.08 1.05
dom-set 18837 96 166 5.78 2.58 41.3 1.30 1.58
vcover 30531 76 110 5.93 2.68 106 1.99 2.19
dense-sub 20021 21 18 8.00 4.79 170 0.60 0.47
The results verify that a step size search strategy significantly reduces the number of MWU iterations compared to the standard step size prescribed in theoretical algorithms. Since an MWU iteration tends to be more expensive than a search step iteration (due to the SpMV), these results suggest that finding accurate step sizes, at the expense of a higher search cost, reduces the overall run time. The performance difference between binary search or Newton’s method is relatively small. While Newton’s method on average requires fewer step size search iterations than binary search, it has more MWU iterations than binary search for the two pure covering problems, dom-set and vcover. The additional MWU iterations observed when using Newton’s method may be attributed to the $(1-\epsilon)$ multiplicative decrease (where $\epsilon=0.1$) applied to step sizes violating the bang-for-buck inequality (16).
7.4. Effect of Software Optimizations
We now analyze the acceleration of an MWU iteration with our software optimizations. To do so, we will compare the execution times of MWU-PETSc and MWU-opt implementations with 68 threads. Later, we will also compare the execution times of MWU-PETSc and MWU-opt implementations for problems with implicit matrix vector multiplication.
7.4.1. Performance Breakdown
First, we consider where the cycles are spent in our MWU-PETSc implementation. Figure 5a shows the fraction of time spent in matrix-vector products (matvec), step size search (search), and other vector operations (vec). The gradients (Line 56) and new direction (Line 7) are included in the vec category while all other vector operations done during step size search are included in the search category.
We observe that both applications and input graphs affect the distribution of execution cycles among these three components. For example, while match, bmatch and dom-set problems spend most of their time during matvec, vcover and dense-sub problems spend more than 50% of their execution time for vec and search operations. matvec takes on average 75%, 78%, and 82% of the execution time for match, bmatch, and dom-set problems, respectively. In contrast, for vcover and dense-sub, matvec takes only 45% and 38% of the execution time on average. Due to this variable behavior, it is crucial to optimize both matrix-vector multiplications and vector operations for MWU.
(a) MWU-PETSc execution time breakdown
(b) OMP Breakdown
(c) Speedups of MWU components with MWU-opt compared to MWU-PETSc
Figure 5. Breakdown of execution times and speedups obtained with MWU-opt for different components.
7.4.2. Shared-Memory Performance Optimizations
Figure 5c shows the speedup obtained by our optimized implementation relative to the PETSc-based implementation when executing on a single node. In the rest of this section, we report geometric mean speedups when referring to average speedup across graphs. For dom-set, the speedup is obtained from using a favorable format (CSB) and minimizing memory accesses. Our optimizations accelerate matvec operations by 1.8x on average. Although vec and search operations also get speedups, their contribution to overall performance is smaller. On the other hand, for match, bmatch, and vcover problems, we can observe the benefit of specialized vertex-incidence matrix vector multiplications. In these cases, matvec operations are 3.28x, 5.06x, and 3.49x faster on average, respectively. For dense-sub problem, we also see benefits of vertex-edge pair matrix and interweaved-identity matrix specializations. matvec operations are 4.64x faster on average.
Moreover, vcover and dense-sub problems spend a large amount of time for vec and search operations. We see that, in these cases, MWU-opt implementation can obtain significant speedups for both vec (6.85x, and 4.08x on average, respectively) and search (8.92x and 5.79x on average, respectively) thanks to fusing and SIMD optimizations.
7.4.3. Distributed-Memory Optimizations
We record runtime improvements in the context of multi-node exeuction MWU-opt over MWU-PETSc in Table 4. In parenthesis is the ratio of matvec product time to matvec communication time for MWU-opt. All experiments are run with 64 OMP threads per MPI process. We observe that for 4 nodes, the use of implicit matrix-vector products accelerates matvec operations in MWU-opt by 1.4-3x compared to MWU-PETSc for all graphs except rgg-24, for which it is slower by 9x. As previously discussed in Section 7.2, MWU-PETSc uses a 1D communication layout, which is more efficient on banded matrices like rgg-24 than our 2D communication layout.
Table 4. Speed-up in runtime of our implicit implementation of the product of the edge-incidence matrix and a vector. The ratio of computation to communication time in MWU-opt is parenthesized.
hollyw orkut rgg-24 kron-21
4 nodes 1.4 (0.57) 3 (1.1) 0.11 (0.09) 3 (0.83)
16 nodes 120 (0.39) 46 (0.19) 0.04 (0.02) 97 (0.18)
64 nodes 186 (0.3) 259 (0.12) 0.09 (0.01) 516 (0.19)
8. Conclusion
Our work demonstrates that approximate positive LP solvers are an efficient and scalable approach for solving a wide range of graph problems. We show that with carefully chosen modifications and implementation of the MWU algorithm from Mahoney et. al. (Mahoney et al., 2016) – namely, a step size search strategy and specialized linear algebra operations that leverage shared and distributed-memory resources – the algorithm exceeds the performance of general purpose LP solvers for finding a $(1+\epsilon)$-relative solution. Our implementation also matches the performance of hand-tuned parallel graph libraries for some graphs.
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0022158.
Appendix A Supplementary Material
A.1. Further Details on Generalized Matching Experiments
Let $G=(V,E)$ be an undirected, unweighted graph. For generalized matching, a vertex $v$ can be matched $b(v)$ times, where $\mathrm{lb}(v)\leq b(v)\leq\mathrm{ub}(v)$ are lower and upper bounds on the number of unique vertices matching with $v$. More precisely, the IP formulation is
(18) $\begin{split}\exists\ \bm{x}\text{ s.t. }&\ \mathrm{lb}(v)\leq\sum\limits_{e\in\mathrm{inc}(v)}x_{e}\leq\mathrm{ub}(v),\forall v\in V,\\ &\ x_{e}\in\{0,1\},\forall e\in E.\end{split}$
Maximum matching is equivalent to generalized matching with $\mathrm{lb}(v)=0$ and $\mathrm{ub}(v)=1,\forall v\in V$, as well as a (maximization) objective function of $\sum\limits_{e}x_{e}$. The LP relaxation is the feasibility mixed packing and covering LP,
$\exists\ \bm{x}\in\mathbf{R}^{m}\text{ s.t. }\bm{Mx}\geq\bm{l},\bm{Mx}\leq\bm{u},\bm{x}\geq\mathbb{0},$
where $\bm{M}$ is the vertex-edge incidence matrix, and $\bm{l},\bm{u}\in\mathbf{R}^{n}$ are the vector of lower and upper bounds for each vertex.
A.2. Dataset Preprocessing
Now, let us describe how to pre-process the Netflix (Bennett et al., 2007) and KDD (Dror et al., 2012) datasets, as detailed in (Makari et al., 2013). Both datasets contain users and items (e.g., movies in Netflix, music tracks in KDD) as vertices, and edges correspond to a user rating an item. This dataset is represented as a bipartite graph, where users and items form the two partitions, and edges go only between vertices in separate partitions. For the number of matchings with each user, we enforce a lower bound of three and upper bound of five. For items, no lower bound is set, but an upper bound of 200 and 2000 is chosen for Netflix and KDD, respectively. Finally, to ensure there is a feasible matching satisfying these bounds, we exclude users with less than ten ratings from the Netflix dataset. After this pre-processing step, the two datasets have 473k and 1.6m vertices, as well as 100m and 252m edges, respectively.
A.3. Gradient Descent with Adaptive Error
Finally, we review the gradient descent algorithm with an adaptive error of (Makari et al., 2013). In short, the algorithm minimizes the convex function via gradient descent,
$\Gamma(\bm{x})=\sum\limits_{i=1}^{m_{P}}y_{i}(\bm{x})+\sum\limits_{i=1}^{m_{C}}z_{i}(\bm{x}),$
where for some $\mu>0$,
$y_{i}(\bm{x})$ $=\mathrm{exp}\left[\mu\cdot(\bm{P}_{i}\bm{x}-1)\right]$
$z_{i}(\bm{x})$ $=\mathrm{exp}\left[\mu\cdot(1-\bm{C}_{i}\bm{x})\right].$
The algorithm contains two error values. There is the error bound $\epsilon$, where the algorithm seeks to find an $\bm{x}$ that is a $(1+\epsilon)$-relative solution. Then there is the internal error bound $\epsilon^{\prime}$, which is used to specify $\mu$ as well as which coordinates of $x_{i}$ to update, and by how much. The authors of (Makari et al., 2013) found that they can set $\epsilon^{\prime}>\epsilon$. For example, when $\epsilon=0.05$, they can choose $\epsilon^{\prime}=1$. Then they run the algorithm until it stagnates, and if $\bm{x}$ is not an $(1+\epsilon)$-relative solution, they decrement $\epsilon^{\prime}$ and warm-start the algorithm by setting the initial $\bm{x}_{0}$ to the solution of the previous, stagnated algorithm. This strategy is called adaptive error, since it adaptively updates the internal error bound.